40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 3 16:43:44 2011, Katrin, Update, Green Locking, SHG temperature (YARM) 
    Reply  Thu Nov 3 17:04:23 2011, Koji, Update, Green Locking, SHG temperature (YARM) 
       Reply  Thu Nov 3 17:24:59 2011, Zach, Update, Green Locking, SHG temperature (YARM) 
          Reply  Fri Nov 4 13:53:33 2011, Katrin, Update, Green Locking, SHG temperature (YARM) 
    Reply  Wed Nov 9 16:49:17 2011, kiwamu, Update, Green Locking, Y end laser temperature with slow input connected 
Message ID: 5852     Entry time: Wed Nov 9 16:49:17 2011     In reply to: 5797
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y end laser temperature with slow input connected 

Indeed it is strange. I took a quick look at it.

In order to recover the same condition (e.g. the same amount of the reflected DC light and the same temperature readout),

it needed to have +8.9V in the slow input from the DAC through EPICS.

Obviously applying an offset in the slow input to maintain the same condition is not good.

It needs another solution to maintain the sweet frequency where the frequency of the PSL and the Y end laser is close in a range of 200 MHz.

Quote from #5797

Plugging in the thermal feedback BNC cable to the laser reduced the DC voltage of the green PDH photo diode from 3.12 V to 1.5V off resonance.

 

ELOG V3.1.3-